OpenStudy (anonymous):

cos(x)-sin(x)=sqrt(2)sin(x) prove that cos(x)+sin(x)=sqrt(2)cos(x)

OpenStudy (alekos):

are you still there?

OpenStudy (alekos):

@RahulYadav-2000

OpenStudy (anonymous):

hey my internet connection is bad that's why it gets disconnected if any body knows the answer pls send it to me

OpenStudy (alekos):

If \[cosx - sinx = \sqrt2 sinx\] then \[(1/\sqrt2) cosx - (1/\sqrt2) sinx = sinx\] and \[\cos \pi/4cosx - \sin \pi/4sinx = sinx\] hence \[\cos (x + \pi/4) = sinx\] so \[\sin(\pi/4 -x) = sinx\] and this leads to \[x=\pi/4\]

OpenStudy (alekos):

sorry it leads to \[x=\pi/8\]

OpenStudy (alekos):

so in order to prove the second statement just substitute \[x = \pi/8\] to both sides

OpenStudy (anonymous):

1.f(x)= log(\[\sqrt{x ^{2}+1}\]-x) 2.f(x)= xlog(\[x+\sqrt{x ^{2} +1}\]) find which is odd and which is even